OpenStudy (anonymous):

The equation of line QR is y = negative 1 over 2x + 1. Write an equation of a line perpendicular to line QR in slope-intercept form that contains point (5, 6). I think its y=2x-4

OpenStudy (anonymous):

@bohotness

OpenStudy (anonymous):

@badmood

OpenStudy (badmood):

Is the equation of Line XY y−3=−2/3(x−4)

OpenStudy (badmood):

sorry wrong thing

OpenStudy (anonymous):

Haha you're good bro

OpenStudy (badmood):

www.jiskha.com/search/index.cgi?query=Find the slope and y...

OpenStudy (badmood):

or homework help.com

OpenStudy (badmood):

1/(2x=1) is not the equation of a line, but of an hyperbola.

OpenStudy (badmood):

got it?